if x^3 + 4x =8 find value of x^7 +64x^2 a)132 b)128 c)216 d)125

if  x^3  + 4x =8 find value of  x^7 +64x^2                         


a)132


b)128 


c)216


d)125

Grade:11

1 Answers

Sher Mohammad IIT Delhi
9 years ago
x^3 +4x=8 square the equation
x^6 +16x^2 +8x^4=64 multiply with x
x^7+16x^3+8x^5-64x=0

now x^7+64x^2 put value of x^7
64x-16x^3-8x^5+64x^2
64x-8x^2(x^3+4x)+16x^3+64x^2
64x-64x^2+16x^3+64x^2
64x+16x^3
16(x^3+4x)=16*8=128
ans is B

Sher Mohammad
faculty askiitians